OpenStudy (anonymous):

Two positive numbers have a difference of 8 and a product of 33. What are these numbers?

OpenStudy (anonymous):

Only whole number factors of 33 are 3 and 11.

OpenStudy (anonymous):

Now where's my medal? XD

OpenStudy (anonymous):

let the first number be x hence other will be x+8 According to the question we have x(x+8) =33 i.e. \[x^2+8x -33= 0\] \[x^2+11x-3x -33= 0 \rightarrow x(x+11) -3(x +11)= 0\] \[ \rightarrow (x+11) (x -3)= 0\] i.e. (x+11) =0 or (x -3)= 0 i.e. x=-11 or x=3 since x is positive therefore, \[x \neq -11\] x= 3 x+8= 3+8= 11 hence the two positive numbers are 3 and 11 VERIFICVATION; Difference: 11-3=8 Product: 11*3 =33 @serrabrana

OpenStudy (anonymous):

@serrabrana

OpenStudy (anonymous):

Where is my medal? XD

OpenStudy (anonymous):

*waits*